An angle measures 155.2° less than the measure of its supplementary angle. What is the measure of each angle?

Answer by greenestamps(13200)



x = measure of the given angle
180-x = measure of its supplementary angle

The angle measure is 155.2 degrees less than the measure of its supplement:

x = (180-x)-155.2
x = 24.8-x
2x = 24.8
x = 12.4

ANSWERS:
angle: 12.4 degrees
supplement: 180-12.4 = 167.6